The length of a rectangle is 4 inches less than twice the width. if the area of the rectangle is 240 sq inches, what is the length and width of the rectangle?
Answer by checkley71(8403)

LENGTH=2W-4
AREA=L*W
240=(2W-4)W
240=2W^2-4W
2W^2-4W-240=0
2(W^2-2W-120)=0
2(W-12)(W+10)=0
W-12=0
W=12 ANSWER FOR THE WIDTH.
THUS THE LENGTH IS:
2*12-4=24-4=20 FOR THE LENGTH.
ROOF
12*20=240
240=240